# Accelerating FRB Search: Dataset and Methods

### v1

@@ Abstract
Fast Radio Burst (FRB) is an extremely energetic cosmic phenomenon of short duration. Discovered only recently and with yet unknown origin, FRBs have already started to play a significant role in studying the distribution and evolution of matter in the universe. FRBs can only be observed through radio telescopes, which produce petabytes of data, rendering the search for FRB a challenging task. Traditional techniques are computationally expensive, time-consuming, and generally biased against weak signals. Various machine learning algorithms have been developed and employed, which all require substantial data sets. We here introduce the FAST dataset for Fast Radio bursts EXploration (FAST-FREX), built upon the observations obtained by the Five-hundred-meter Aperture Spherical radio Telescope (FAST). Our dataset comprises 600 positive samples of observed FRB signals from three sources and 1000 negative samples of noise and Radio Frequency Interference (RFI). Furthermore, we provide a machine learning algorithm, Radio Single-Pulse Detection Algorithm Based on Visual Morphological Features (RaSPDAM), with significant improvements in efficiency and accuracy for FRB search. We also employed the benchmark comparison between conventional single-pulse search softwares, namely PRESTO and Heimdall, and RaSPDAM. Future machine learning algorithms can use this as a reference point to measure their performance and help the potential improvements.
@@ 1 Introduction
Fast Radio Burst (FRB) is a rapidly developing field in astrophysics that researchers extensively explore regarding its origin, emission mechanisms, and applications. FRBs are characterized by their short duration, high-energy release, and widespread distribution in the universe, which is crucial in studying the distribution and evolution of matter in the cosmos. Despite the discovery of about a few hundred FRBs (Li et al., 2021; Niu et al., 2022a; Feng et al., 2022; Zhang et al., 2022), their origin and physical mechanisms are still a mystery. The discovery of more FRBs is expected to bring revolutionary breakthroughs in physics and astronomy.
Various open-source software packages have been developed for searching single-pulse signals to discover FRBs, long-period pulsars, and other radio transient sources. PRESTO (Ransom, 2011) and SIGPROC (Lorimer, 2011) are two software packages designed to standardize the initial analysis of various fast-sampled pulsar data types and have aided in the discovery of numerous pulsar radio sources. Heimdall  [^1] https://sourceforge.net/p/heimdall-astro/wiki/Home/, a single-pulse search algorithm based on Graphical Processing Units (GPUs), can enhance data processing efficiency to a certain degree. These software packages’ pipelines usually employ computation-intensive dispersion removal algorithms on search data, producing several time sequences with dispersion effects removed, which might lead to significant computational time. On the other hand, Radio Frequency Interference (RFI) and instrument noise make detecting weaker pulse signals challenging.
@@
Astronomical telescopes have gathered extensive observational data that aids in the search for FRBs. The Chinese Five-hundred-meter Aperture Spherical radio Telescope (FAST) (Li et al., 2018; Jiang et al., 2019), which is the world’s biggest single-aperture telescope with the highest sensitivity, can detect weak signals that other radio telescopes cannot detect due to its precise surface reflection control and the favorable polarization characteristics of its FAST-19 beam receiver. The research team utilized FAST to detect 1652 outbursts in about a 50-day interval through the Commensal Radio Astronomy FAST Survey (CRAFTS), obtaining the most extensive sample of FRB outbursts to date and revealing the complete spectrum and bimodal structure of the burst rate for the first time (Li et al., 2021).
@@ 3.1 Data Collection
Our objective is to develop and compare algorithms that can identify FRB events. Therefore, we selected 600 bursts to create positive data samples and ensured that each positive sample contained only one FRB event. In order to preserve the integrity of FRB events, we set the cropping duration of each positive sample file to about 6 seconds. We also ensure that the events appear randomly within the observation time covered by the file rather than setting a fixed Time of Arrival (ToA). This approach simulates FRB detection more realistically and enhances the dataset’s diversity, which is beneficial for algorithm research. We reduced the data size by averaging the first two polarization signals while maintaining the signal’s integrity and mitigating the noise’s effect to some extent. After eliminating the records of authentic events, we extracted 1000 negative samples of RFI and noise with the same duration as the positive samples.
@@ 3.2.1 Authentic Single Pulses
The FWHM is a parameter commonly used to describe the width of the burst profile. It is given by the distance between points on the spectral curve at which the intensity reaches half its peak. The burst is distributed over a frequency interval, and the range of this frequency interval is called bandwidth. The spectral flux density is a quantity that indicates the observed strength of an astronomical source. It is a measure of the strength of a radio signal received from a discrete source, which is measured in watts per square meter per hertz ($\rm{W~{}m^{-2}~{}Hz^{-1}}$). Peak flux density $S_{\rm{peak}}$ is the maximum flux intensity of the pulse profile. The unit of flux density is the jansky (Jy), and 1 Jy equals $\rm{10^{-26}~{}W~{}m^{-2}~{}Hz^{-1}}$. The fluence F of a burst is computed by integrating the burst profile concerning time and bandwidth, which is used to describe the total energy of the burst. The unit of F is Jy ms.
@@
In Figure 2, we present the frequency-time domain representation of FRB signals, varying pulse width of positive samples. Pulse width indicates the duration of the FRB signals on the time axis. Figure 3 displays the frequency-time domain of FRB signals, with varying signal flux densities of positive samples. This visualization illustrates that signals with higher flux density are more prominently observed than the background noise.
Figure 2: Examples of positive samples with different pulse widths
Figure 3: Examples of positive samples with different flux densities
@@ 4.1 RaSPDAM
In conventional approaches, the initial step of searching for FRBs involves de-dispersion, a process that requires significant time and computational resources (CPU/GPU cores) to perform the Fast Fourier Transform (FFT) operation. As a result, the process of searching for FRBs has become computationally intensive and time-consuming. In the context of RaSPDAM, observational data is processed as conventional images. The signal sequence is segmented and converted into an image format with dimensions of $512\times 512$ pixels. The image serves as the input for a semantic segmentation model.
@@ 4.1.1 Signal Preprocessing & Enhancement
The initial step is to convert the original signal sequence into standard images without de-dispersion. The signal sequence is divided into smaller slices within a 2-second width sliding window and a 1-second overlap. To standardize the model input, we resized the image slices to dimensions of $512\times 512$.
@@ 4.1.2 Model Training
The aim of FRB detection resembles semantic segmentation, also known as pixel-based classification. It requires a method to differentiate the FRB signal component from the background, where noise might interfere with segmentation outcomes. In this case, we choose the UNet as the model implementation due to several key advantages:
@@
Unet is a Convolutional Neural Network specialized in classification tasks, localizing specific areas by classifying individual pixels. The architecture resembles a U-shape, with the left side part known as the encoder or contracting path and the right side as the decoder or expansion path.
Training the network requires a significant amount of FRB signal images. However, there is a lack of available training data in the field of astronomy. Our proposed alternative involves generating simulated FRB signals. As previously mentioned, we enhance the signal sequence slice using convolution kernels. To generate simulated signals, we apply a similar process by using random curve slopes and Gaussian noises. With the known slope of the FRB curve, it is possible to accurately determine its position and create mask images for training purposes.
@@ 4.1.3 Candidate Identification
The network segmentation process produces segmentation maps for the purpose of identifying potential candidates for FRB signals. However, it may still contain noise. In response to this issue, we have implemented multiple procedures for candidate identification.
Figure 6 shows that we utilize the ”regionprops” function for the analysis of connected areas. Each identified candidate is quantified as x, while the area of the region box and the filled area are represented as area_bbox and area_filled. Then, we use thresholds t_a. If t_a is smaller than the ratio of area_filled to area_bbox, it is flagged as noise, as indicated in Equation 2.
@@
Additionally, we calculate the candidate’s projection on both the x-axis and y-axis. Candidates exceeding both x-axis and y-axis projections above thresholds that denoted as t_b and t_c, as potential FRB signals.
@@ 5.1 Baseline Scoring Metrics
We aim to improve the efficiency and accuracy of identifying positive samples containing FRB signals. Consequently, our approach yields metrics, including recall and precision rates. The True Positives (TPs) and True Negatives (TNs) mean the correctly predicted positive candidates and negative samples. Meanwhile, False Positives (FPs) and False Negatives (FNs) reflect the erroneously predicted positive candidates and negative samples. We define that detecting FRB signal candidates in positive samples with a ToA error $\leq$ 0.2s and a DM error $\leq$ 50 pc $cm^{-3}$ are correctly predicted TPs. Candidates in positive samples that do not meet the specified error thresholds and candidates found in negative samples are FPs. Consequently, the number of FPs may exceed the number of sample files, as each sample file could yield multiple candidates. If no FRB signal candidates appear in negative samples, they are correctly predicted as negative samples (TNs). If no candidates are found in positive samples, it is treated as FNs.
@@ 5.2 Results
RaSPDAM 1000 600 1600 989 466 128 6 0.7767 0.9873 0.8694
The results indicate that PRESTO and Heimdall have a higher recall rate than the machine learning algorithm RaSPDAM on the total dataset. However, the accuracy of RaSPDAM is very high, whereas both PRESTO and Heimdall are unsatisfactory. Due to the misidentification of RFI as genuine signals, PRESTO and Heimdall calculate a significantly larger number of detected signals than true FRBs. As a result, the absence of human intervention makes it difficult for the two conventional softwares to screen out true FRB signals effectively.
@@ 6.1 Distributed Computing Architecture
To address the challenge of processing petabyte-scale data generated by FAST, we adopted a distributed computing architecture based on Kubernetes  [^4] https://kubernetes.io/ and Airflow  [^5] https://airflow.apache.org/. This architecture significantly extends our data processing capabilities, enabling the efficient handling of massive datasets.
@@ 6.2 Discoveries of RaSPDAM
The RaSPDAM algorithm has been successfully implemented in FRB and pulsar studies. Since its deployment last year, our team has leveraged this algorithm to identify 2 FRBs (FRB20211103A and FRB20230104) and 80 pulsars. Among these pulsar detections, 13 are previously undiscovered pulsars, highlighting the algorithm’s efficacy in uncovering new celestial objects. Table 3 presents a compilation of these pulsars.
@@ 7 Discussion
2) Our dataset currently only includes 600 signals from three FRB sources and RFI/noise observed from the specific observations. However, due to variations in observational conditions and intrinsic properties of FRBs, unexpected patterns possibly exist in unknown candidates, which causes models trained on this dataset to fail to recognize other signals. The Blinkverse (Xu et al., 2023) database accumulated an impressive catalog of 8,007 FRB bursts from 813 FRB sources (data up to May 27, 2024), including the bursts our dataset collected. In future work, we aim to extend the dataset with more FRBs, guided by the comprehensive records within Blinkverse, to enhance the comprehensiveness of our dataset.
Furthermore, RaSPDAM can only provide the ToA as a result, while other algorithms provide both ToA and DM, which may cause inconvenience in candidate verification. For future work, we will continue to enhance the capabilities of RaSPDAM to provide more comprehensive features. In addition, in previous model training, we only used simulated FRB signals, which may differ from real signals. We will try to fine-tune the model with real signals in the future, which may further improve the recall rate.

### v2

@@ Abstract
Fast Radio Burst (FRB) is an extremely energetic cosmic phenomenon of short duration. Discovered only recently and with its origin still unknown, FRBs have already started to play a significant role in studying the distribution and evolution of matter in the universe. FRBs can only be observed through radio telescopes, which produce petabytes of data, rendering the search for FRB a challenging task. Traditional techniques are computationally expensive, time-consuming, and generally biased against weak signals. Various machine learning algorithms have been developed and employed, all of which require substantial datasets. We here introduce the FAST dataset for Fast Radio bursts EXploration (FAST-FREX), built upon the observations obtained by the Five-hundred-meter Aperture Spherical radio Telescope (FAST). Our dataset comprises 600 positive samples of observed FRB signals from three sources and 1000 negative samples of noise and Radio Frequency Interference (RFI). Furthermore, we provide a machine learning algorithm, Radio Single-Pulse Detection Algorithm Based on Visual Morphological Features (RaSPDAM), with significant improvements in efficiency and accuracy for FRB search. We also employed the benchmark comparison between conventional single-pulse search softwares, namely PRESTO and Heimdall, and RaSPDAM. RaSPDAMv2 achieves an average precision of 97% and an average recall of 83%, with notable enhancements in computational performance. Future machine learning algorithms can use this as a reference point to measure their performance and help the potential improvements. By enabling more accurate and efficient detection of transient radio events, our work facilitates the FRB and pulsars search pipeline, enhances the potential for discovering new astrophysical phenomena.
@@ 1 Introduction
Fast Radio Burst (FRB) represents a rapidly developing field in astrophysics, with extensive research focusing on its origin, emission mechanisms, and applications. FRBs are characterized by their short duration, high-energy release, and widespread distribution across the universe, which plays a crucial role in studying the distribution and evolution of matter in the cosmos. Despite the discovery of about a few hundred FRBs (Li et al., 2021; Niu et al., 2022a; Feng et al., 2022; Zhang et al., 2022), their origin and physical mechanisms are still a mystery. The discovery of more FRBs is expected to bring revolutionary breakthroughs in physics and astronomy.
Various open-source software packages have been developed for searching single-pulse signals to discover FRBs, long-period pulsars, and other radio transient sources. PRESTO (Ransom, 2011) and SIGPROC (Lorimer, 2011) are two software packages designed to standardize the initial analysis of various fast-sampled pulsar data types and have aided in the discovery of numerous pulsar radio sources. Heimdall  [^1] https://sourceforge.net/p/heimdall-astro/wiki/Home/, a single-pulse search algorithm based on Graphical Processing Units (GPUs), can enhance data processing efficiency to a certain degree. The pipelines of these software packages typically employ computation-intensive dispersion removal algorithms on search data, producing several time sequences with dispersion effects removed, which might lead to significant computational time. On the other hand, Radio Frequency Interference (RFI) and instrument noise make detecting weaker pulse signals challenging.
@@
Astronomical telescopes have gathered extensive observational data that aids in the search for FRBs. The Chinese Five-hundred-meter Aperture Spherical radio Telescope (FAST) (Nan et al., 2011; Li et al., 2018), which is the world’s biggest single-aperture telescope with the highest sensitivity, can detect weak signals that other radio telescopes cannot detect due to its precise surface reflection control and the favorable polarization characteristics of its FAST-19 beam receiver. The research team utilized FAST to detect 1652 outbursts in about a 50-day interval through the Commensal Radio Astronomy FAST Survey (CRAFTS), obtaining the most extensive sample of FRB outbursts to date and revealing the complete spectrum and bimodal structure of the burst rate for the first time (Li et al., 2021).
@@ 3.1 Data Collection
Our objective is to develop and compare algorithms that can identify FRB events. Therefore, we selected 600 bursts to create positive samples and ensured that each positive sample contained only one FRB event. Positive samples refer to data instances that contain confirmed FRB signals. In order to preserve the integrity of FRB events, we set the cropping duration of each positive sample file to about 6 seconds. We also ensure that the events appear randomly within the observation time covered by the file rather than setting a fixed Time of Arrival (ToA). This approach simulates FRB detection more realistically and enhances the dataset’s diversity. We reduced the data size by averaging the first two polarization signals while maintaining the signal’s integrity and mitigating the noise’s effect to some extent. After eliminating the records of authentic events, we extracted 1000 negative samples of RFI and noise with the same duration as the positive samples. Negative samples consist of data instances that do not contain FRB signals.
@@ 3.2.1 Authentic Single Pulses
The FWHM is a parameter commonly used to describe the width of the burst profile. It is given by the distance between points on the spectral curve at which the intensity reaches half its peak. The burst is distributed over a frequency interval, and the range of this frequency interval is called bandwidth. The spectral flux density is a quantity that indicates the observed strength of an astronomical source. It is a measure of the strength of a radio signal received from a discrete source, which is measured in watts per square meter per hertz ($\rm{W~m^{-2}~Hz^{-1}}$). Peak flux density $S_{\rm{peak}}$ is the maximum flux intensity of the pulse profile. The unit of flux density is the jansky (Jy), and 1 Jy equals $\rm{10^{-26}~W~m^{-2}~Hz^{-1}}$. The fluence F of a burst is computed by integrating the burst profile concerning time and bandwidth, which is used to describe the total energy of the burst. The unit of F is Jy ms. In Appendix 4, we provide a detailed presentation of the parameter information for the positive samples in the dataset.
@@
In Figure 2, we present the frequency-time domain representation of FRB signals, varying pulse width of positive samples. Figure 3 displays the frequency-time domain of FRB signals, with varying signal flux densities of positive samples. These visualizations help illustrate how signal characteristics affect detectability.
Figure 2: Examples of positive samples with different pulse widths. Each subplot shows the frequency-time domain (dynamic spectrum) of an FRB signal with a different pulse width. The black solid line represents the frequency-averaged Stokes I time series profile. Each Gaussian pulse component is convolved with the same scattering exponent, and the model fit is shown as a bold red line.
Figure 3: Examples of positive samples with different flux densities. Each subplot displays the frequency-time domain (dynamic spectrum) of an FRB signal with a different flux density. The black solid line indicates the frequency-averaged Stokes I time series profile, while the bold red line represents the model fit. Signals with higher flux densities are more clearly distinguishable from background noise.
@@ 4.1 RaSPDAM
In conventional approaches, the initial step of searching for FRBs involves de-dispersion, a process that requires significant time and computational resources (CPU/GPU cores) to perform the Fast Fourier Transform (FFT) operation. As a result, the process of searching for FRBs has become computationally intensive and time-consuming. In the context of RaSPDAM, observational data is processed as conventional images. The signal sequence is segmented and converted into an image format with dimensions of $\times 512$ 512 pixels. The image serves as the input for a semantic segmentation model.
@@ 4.1.1 Signal Preprocessing & Enhancement
The initial step is to convert the original signal sequence into standard images without de-dispersion. The signal sequence is divided into smaller slices within a 2-second width sliding window and a 1-second overlap to control temporal resolution. To standardize the model input, we resized the image slices to dimensions of $\times 512$ 512.
@@ 4.1.2 Model Training
The aim of FRB detection resembles semantic segmentation, also known as pixel-based classification, which facilitates subsequent filtering and post-processing steps. It requires a method to differentiate the FRB signal component from the background, where noise might interfere with segmentation outcomes. Models like ResNet (He et al., 2015) or EfficientNet (Tan & Le, 2019) can serve as encoders for semantic segmentation. However, they lack dedicated decoder structures and the critical skip connection mechanism needed to restore spatial details and generate pixel-level segmentation maps. In contrast, UNet(Ronneberger et al., 2015) is a complete end-to-end CNN-based segmentation architecture. It explicitly features a symmetric encoder-decoder structure with skip connections, making it exceptionally suited for fine-grained boundary segmentation tasks. Given these advantages, we adopt the UNet architecture as backbone of RaSPDAM to segment time-frequency spectrograms. Its key benefits include:
@@
RaSPDAMv1 was built upon a standard UNet model, trained using RMSprop (learning rate: 5e-5, momentum: 0.999, weight decay: 1e-8) to ensure stable convergence. The loss function combined Cross-Entropy and Dice Loss, balancing pixel-level accuracy and structural consistency. This served as a solid baseline for further development in RaSPDAMv2.
During the training of RaSPDAMv2, we employed a self-adapting framework, nnUNet (Isensee et al., 2020), which also incorporates conceptual advancements from ResNet, to train our model. We opt for 5-fold training, and this framework employs a 1000-epoch schedule for high-resolution 2D medical image segmentation. Training uses full $\times 512$ 512 patches (batch size=12) without sub-volume sampling, thus preserving spatial context. It is initialized with a base learning rate of 0.01, and a polynomial decay strategy gradually reduces the learning rate to ensure stable convergence. The batch Dice loss addresses class imbalance by evaluating batch-level segmentation consistency rather than focusing on individual samples. Data normalization is performed using Z-score normalization across all three channels, utilizing precomputed statistics. Input images are resampled using cubic interpolation, while segmentation masks are resampled using linear interpolation to maintain anatomical boundaries.
Training the network requires a significant amount of FRB signal images. However, there is a lack of available training data in the field of astronomy. Our proposed alternative involves generating simulated FRB signals. As mentioned above, we enhance the signal sequence slice using convolution kernels. To generate simulated signals, we apply a similar process by using random curve slopes and Gaussian noises. With the known slope of the FRB curve, it is possible to accurately determine its position and create mask images for training purposes.
@@ 4.1.3 Candidate Identification
The network segmentation process produces segmentation maps for the purpose of identifying potential candidates for FRB signals. To eliminate duplicate candidates generated during the network segmentation process, we implemented a two-stage filtering strategy:
  1. 1.
    Non-Maximum Suppression (NMS) with an IoU threshold of 0.5 (default) to remove overlapping bounding boxes;
  2. 2.
    An additional overlap analysis using a containment threshold of 0.5 (default), specifically designed to eliminate boxes that are largely enclosed by others.
However, it may still contain noise. In response to this issue, we have implemented multiple procedures for candidate identification.
Figure 6 shows that we utilize the ”regionprops” function for the analysis of connected areas. Each identified candidate is quantified as x, while the area of the region box and the filled area are represented as area_bbox and area_filled. Then, we use thresholds t_a (default: 0.4). If t_a is smaller than the ratio of area_filled to area_bbox, it is flagged as noise, as indicated in Equation 2.
@@
Additionally, we calculate the candidate’s projection on both the x-axis and y-axis. Candidates exceeding both x-axis and y-axis projections above thresholds (default: 0.05) that denoted as t_b and t_c, as potential FRB signals.
@@ 5.1 Baseline Scoring Metrics
We aim to improve the efficiency and accuracy of identifying positive samples containing FRB signals. Consequently, our approach yields metrics, including recall and precision rates. The True Positives (TPs) and True Negatives (TNs) mean the correctly predicted positive candidates and negative samples. Meanwhile, False Positives (FPs) and False Negatives (FNs) reflect the erroneously predicted positive candidates and negative samples. We define that detecting FRB signal candidates in positive samples with a ToA error $\leq$ 0.2s and a DM error $\leq$ 50 pc $cm^{-3}$ are correctly predicted TPs. Candidates in positive samples that do not meet the specified error thresholds and candidates found in negative samples are FPs. Consequently, the number of FPs may exceed the number of sample files, as each sample file could yield multiple candidates. If no FRB signal candidates appear in negative samples, they are correctly predicted as negative samples (TNs). If no candidates are found in positive samples, it is treated as FNs. In real-world FRB detection scenarios, FPs usually rise from noise or RFI. FNs usually occur when the signal is too weak to be detected by the algorithm, which can happen due to various factors such as low SNR, instrumental limitations, or algorithmic thresholds. FN examples are shown in Figure 7.
Figure 7: FN examples
@@ 5.2 Results
RaSPDAMv1(UNet) 1000 600 1600 989 466 128 6 0.7767 0.9873 0.8694
RaSPDAMv2(nnUNet) 1000 600 1600 994 501 67 14 0.8350 0.9728 0.9253
The results show that PRESTO and Heimdall achieve higher recall rates than the machine learning algorithm RaSPDAMv1 on the dataset. However, RaSPDAMv1 demonstrates superior precision, whereas both PRESTO and Heimdall exhibit unsatisfactory performance in this regard. RaSPDAMv2, on the other hand, significantly improves both precision and recall. Due to the misidentification of RFI as genuine signals, PRESTO and Heimdall calculate a significantly larger number of detected signals than true FRBs. As a result, the absence of human intervention makes it difficult for the two conventional softwares to screen out true FRB signals effectively.
@@ 6.1 Distributed Computing Architecture
To address the challenge of processing petabyte-scale data generated by FAST, we adopted a distributed computing architecture based on Kubernetes  [^4] https://kubernetes.io/ and Airflow  [^5] https://airflow.apache.org/, shown in Figure 9. This architecture significantly extends our data processing capabilities, enabling efficient handling of massive datasets.
Figure 9: Distributed Architecture of Data Processing
@@ 6.2 Discoveries of RaSPDAM
The RaSPDAM algorithm has been successfully implemented in FRB and pulsar studies. Since its deployment last year, our team has leveraged this algorithm to process nearly 8PB observational data from the FAST CRAFTS project, and identified 2 FRBs (FRB20211103A and FRB20230104) and 77 pulsars. Among these pulsar detections, 13 are previously undiscovered pulsars, highlighting the algorithm’s efficacy in uncovering new celestial objects. Table 3 presents a compilation of these pulsars.
@@ 7 Discussion
2) Our dataset currently only includes 600 signals from three FRB sources and RFI/noise observed from the specific observations. However, due to variations in observational conditions and intrinsic properties of FRBs, unexpected patterns possibly exist in unknown candidates, which causes models trained on this dataset to fail to recognize other signals. The Blinkverse (Xu et al., 2023) database accumulated an impressive catalog of 8,007 FRB bursts from 813 FRB sources (data up to May 27, 2024), including the bursts our dataset collected. In future work, we aim to extend the dataset with more FRBs, guided by the comprehensive records within Blinkverse, to enhance the comprehensiveness of our dataset. Additionally, we plan to incorporate simulated data into the dataset and will release the associated simulation signal generation algorithm as open-source software.
RaSPDAM demonstrates promising performance in FRB detection. However, several limitations remain to be addressed in future work:
1) Currently, RaSPDAM only outputs the ToA of detected signals, whereas traditional algorithms such as PRESTO and Heimdall also provide DM estimates. The absence of DM may introduce inconvenience during candidate verification. Future enhancements will focus on extending RaSPDAM’s capabilities to include both ToA and DM estimation for more comprehensive results.
2) In its current form, RaSPDAM does not exploit discontinuous signals within the same burst when there are multiple sources of interference. This is an important direction for future work, as leveraging these discontinuities could potentially improve detection accuracy and signal interpretation in more complex scenarios.
3) The current version of RaSPDAM struggles to detect faint or weak signals, which is a critical limitation in practical applications. Further research is required to enhance the algorithm’s sensitivity, particularly in low signal-to-noise ratio (SNR) scenarios. We also plan to explore Transformer-based extensions of the UNet architecture like SwinUnet(Cao et al., 2021) to enhance the model’s capability in detecting weak and low-SNR FRB signals.
4) The model was trained exclusively on simulated FRB signals, which may differ from real-world data in terms of morphology and noise characteristics. Future work will involve fine-tuning the model using real observed signals to improve generalization and recall rate.
5) While RaSPDAM has been optimized for use with FAST data, its applicability to other telescopes remains untested. We plan to adapt the algorithm for use with data from other observatories, including GBT, to ensure seamless integration into diverse FRB search pipelines. This adaptation is essential for validating RaSPDAM’s performance under diverse data characteristics and observational conditions.